# Entity-based Reinforcement Learning for Autonomous Cyber Defence

### v2

@@ 1 Introduction
The development of reliable agents for autonomous cyber defence is a growing research area aimed at addressing the increasing complexity and volume of cyber threats. Solutions based on deep reinforcement learning (RL), inspired by high-profile successes in other complex domains (Vinyals et al.,, 2019; Roy et al.,, 2021; Bello et al.,, 2017), offer the potential to rapidly detect, analyse, and respond to cyber attacks at scale (Nguyen and Reddi,, 2021). A fundamental challenge involved in the deep RL approach is to develop agents that are generalisable to variable network topologies. This is difficult to achieve with fixed-input neural network architectures commonly used for function approximation and standard RL interfaces such as Gym or Gymnasium (Brockman et al.,, 2016; Towers et al.,, 2024) which expect fixed-size observation vectors and action spaces. A lack of this generalisation ability imposes serious limitations on the potential deployment of RL agents in real-world network scenarios, as enterprise networks can vary dramatically in size, structure, and complexity.
@@
This problem falls under the broad category of compositional generalisation, which is a long-studied problem in reinforcement learning (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Diuk et al.,, 2008). Recently, a number of works have explored introducing compositional generalisation to deep reinforcement learning  (Lin et al., 2023a, ; Mambelli et al.,, 2022; Zhou et al.,, 2022; Haramati et al.,, 2024), through using permutation invariant policy architectures based on attention (Vaswani et al.,, 2017) or Deep Sets (Zaheer et al.,, 2017). Indeed, some existing works relating to autonomous cyber defence have suggested using Transformers to parameterise a value function (Mern et al.,, 2021), or using graph neural networks for policies (Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023) which are invariant to the number of nodes in the environment.
@@
This work aims to address this challenge in the context of autonomous cyber defence by leveraging entity-based RL (Winter et al., 2023a, ). Entity-based RL is a paradigm where environments are decomposed into collections of discrete entities which an agent observes and acts upon directly. In the context of a network environment, each node of the network can be treated as an entity, with a defending agent’s observation space permitted to vary between environment instances according to how many nodes are visible. With a suitable training regime, it should be possible for an agent trained using entity-based RL to generalise to any network topology, provided the network in question contains familiar types of entities. The Entity-Gym framework (Winter et al., 2023b, ), intended to occupy a similar role as OpenAI Gym or Gymnasium, provides a standard interface in which to cast entity-based environments, and allows any agent designed for entity-based reinforcement learning to be seamlessly deployed on such environments.
@@
In this paper, we introduce a wrapper for the Yawning Titan cyber simulation environment (Andrew et al.,, 2022) that adheres to the Entity Gym interface. On this environment, we trained and compared two types of policies: a Transformer-based policy, RogueNet (Winter et al., 2023a, ), trained on the entity-based interface and a multilayer perceptron (MLP) policy trained on the standard Gym (Brockman et al.,, 2016) interface for Yawning Titan. We trained both types of policies using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017) with two different training modes, which we refer to as Random and Static, across network sizes of 10, 20, and 40 nodes. In the Random mode, a new random network topology is generated with each environment reset, whilst in the Static mode, a random network topology is initialised at the start of training and kept constant throughout. This approach has been designed to facilitate comparison with the MLP-based policy, which is able to train well in the Static mode, but struggles with Random training where it encounters a new network configuration with every episode. We also performed zero-shot evaluation of the Transformer-based policies on network sizes not seen during training, something not possible with standard MLP-based policies. From these experiments we find that in this environment, entity-based policies outperform MLP-based policies during training, particularly in the Random mode where the MLP-based policy struggled to learn. We also find that entity-based policies generalise well to networks of unseen sizes, performing similarly to policies trained natively in those environments.
@@
The paper is organised as follows. In section 2 we give an overview of motivating background and related literature. In section 3 we describe the Entity Gym and the RogueNet architecture (Winter et al., 2023a, ; Winter et al., 2023b, ) which we have used to construct our environment and train entity-based policies. In section 4 we provide a description of the Yawning Titan simulator, and choices we have made for the environment setting for our experiments. In section 5 we explain how we have adapted the Yawning Titan simulator to the entity-based framework. In section 6 we outline the methodology and experiments we have chosen to run. In section 7, we present the results of our experiments and conclude with a discussion of results and future work in section 8.
@@ 2.1 Autonomous Cyber Defence
The CybORG environment (Standen et al., 2022b, ) was initially conceived as part of a competition to encourage research into RL-based autonomous cyber defence agents, Cage Challenge 1 (Standen et al., 2021a, ). High scoring solutions to the challenge used agents trained with PPO (Schulman et al.,, 2017), combined with a bandit controller to select between policies specialised in specific attack patterns (Foley et al., 2022a, ; Hicks et al.,, 2023; Foley et al., 2022b, ; Bates et al.,, 2023), or with an ensemble approach (Wolk et al.,, 2022). Whilst not the focus of our paper, the CybORG environment features several types of nodes that could be used as entity types in the entity-based RL setting; User Hosts, Enterprise Servers, Operational Hosts and the Operational Server.
Recognising that training and evaluating on a purely fixed network environment is not satisfactory for real-world applicability, a number of works have explored forms of generalisation. For example, to unseen red agent dynamics (Wolk et al.,, 2022), or to unseen network topologies, whether that be on fixed networks with variable connectivity (Collyer et al.,, 2022), or to networks containing differing numbers of nodes (Applebaum et al.,, 2022; Mern et al.,, 2021; Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023). Indeed, the recent CAGE Challenge 4 (TTCP CAGE Working Group,, 2023) introduced a level of variability in the number of possible user nodes to encourage solutions addressing this problem.
@@
To the best of our knowledge, the most relevant prior work is Mern et al., (2021, 2022), where the authors implement a reinforcement learning environment for the autonomous defence of an industrial control network, and use an attention-based action-value function. Similarly to the ‘entity-based’ paradigm, they use separate modules to encode different node-types, and share encoder parameters between all nodes of the same type. A global attention layer is then used to pass information between all node embeddings, before action-value estimates are decoded for all actions on all nodes. The parameters are learnt using a version of the Deep Q-Learning algorithm (Mnih et al.,, 2013; Hessel et al.,, 2017). This differs from the explicit policy parameterisation used in this work, trained using PPO. An advantage of a policy-based approach is that it allows for a more natural implementation of composite action spaces.
@@ 2.2 Entities and Objects in Reinforcement Learning
In this section we provide an overview of attempts to leverage the underlying structure of RL environments, as well as practical approaches for dealing with complex environments with potentially variable observation and action spaces, such as in the case of autonomous cyber defence. In the first instance, particularly in the case of Factored MDPs (Boutilier et al.,, 2000), the motivation is to find more efficient solution methods by exploiting the ‘factorised’ structure of an environment. More broadly, all of these approaches contribute to solving the problem of compositional generalisation (Zhao et al.,, 2022; Lin et al., 2023b, ). That is, training policies that are able to generalise across environments containing different compositions of familiar objects or entities and dynamics. For further reference, Mohan et al., (2024) provide a comprehensive survey of structure in reinforcement learning.
@@ 2.2.1 AlphaStar and entity-based Reinforcement learning
AlphaStar (Vinyals et al.,, 2019; Choi,, 2020) is an agent developed to play the game Starcraft II at human professional level. Starcraft II is a real-time strategy game where two opposing players develop a base of operations and direct armies of units with the goal of defeating their opponent. Starcraft is a hard problem, both in terms of the practicalities of agent design and the complexity of strategies, which captures many of the challenges of real-world deployment of autonomous agents. Of particular relevance is that the number and nature of the entities an agent can observe and act upon varies throughout the game. This problem may be similar to that of nodes joining or leaving a network, or more generally the problem of transferring to an unfamiliar network consisting of familiar entities. Rather than using purely image-based inputs of the game screen as observations, AlphaStar maintains an entity list consisting of every visible object the agent can interact with. This could be army units, buildings, resources, or terrain. Crucially, this list will vary throughout a game, and between games.
Each entity is associated with a one-dimensional vector that summarises the state and properties of the entity. AlphaStar encodes these entity observations through a three-layer Transformer (Vaswani et al.,, 2017) to learn the relationships between entities. The entity embeddings generated by the Transformer are merged with other scalar and game-map information to form a joint representation for a single time-step’s observation, which is used as input to an LSTM (Hochreiter and Schmidhuber,, 1997) that learns temporal dependencies.
@@ 2.2.2 Factored MDPs
Factored MDPs (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Osband and Van Roy,, 2014) provide a framework that allows for more efficient learning on environments that may be naturally ‘factorised’ into distinct components. Here, the reward and transition functions can also be factorised according to these components, and display conditional independence properties. Guestrin et al., (2003) introduced efficient solution methods for reinforcement learning on factored MDPs as well as a toy environment, SysAdmin, on which to study them. In this environment, a system administrator must maintain a network of computers, each of which can be functional or failed. The administrator must decide whether to reboot one of the machines or do nothing at each time step, aiming to maximise rewards received for working machines, while considering that machine failures are influenced by the state of neighbouring machines. In this environment, the transition of each individual machine is conditionally independent of all other machines given its immediate neighbours, hence it can be cast as a factored MDP. The SysAdmin example instinctively bears resemblance to autonomous cyber defence environments, and F-MDP framework can be a suitable candidate to express network security problems such as the ones considered in this work.
@@ 2.2.3 Relational Reinforcement Learning
Zambaldi et al., (2019) study ‘deep’ relational reinforcement learning, and advocate the use of GNN or attention-based function approximators in order to learn abstract representations of relations between entities in an environment. This is opposed to traditional relational RL, where relations are provided a-priori. They find the use of ‘relational’ attention modules allows an agent to generalise zero-shot to unseen environment configurations. This approach essentially discards the traditional formal solution methods based on logical descriptions of the environment. In other words, it is a purely ‘entity-based’ approach, where the environment is split into distinct entities, and any relations between them are learnt during training.
@@ 2.2.5 Graph-based Reinforcement Learning for robotic control
A number of papers have studied the use of GNN or Transformer-based policies on robotics continuous control environments, which Hong et al., (2022) refers to as ‘Inhomogeneous Multitask Reinforcement Learning’. Here, the agent is tasked with learning policies across multiple tasks that may have different (inhomogeneous) state and action spaces or dynamics. The goal is to leverage shared knowledge and experience across these tasks to improve learning efficiency and performance. This has been successfully applied to robotics environments (Wang et al.,, 2018; Huang et al.,, 2020; Kurin et al.,, 2021; Hong et al.,, 2022), where the observation and action space are taken to be nodes in a graph of joints of the robot. An agent may be trained across different robot morphologies given by different graphs, with the goal that the policy is generalisable across all of these settings. Such tasks can be formalised by considering a set of MDPs $\mathcal{M}^{K}=\{M_{1},\dots,M_{K}\}$ (Kurin et al.,, 2021; Hong et al.,, 2022). The set $\mathcal{M}^{K}$ is inhomogeneous if there exists $i,j\in[K],i\neq j$ such that $dim(\mathcal{S}_{i})\neq dim(\mathcal{S}_{j})$ or $dim(\mathcal{A}_{i})\neq dim(\mathcal{A}_{j})$, where $\mathcal{S}_{i}$ and $\mathcal{A}_{i}$ refer to the state and action spaces of MDP $M_{i}$. The aim is then to construct a policy that maximises the average expected discounted return over all environments in $\mathcal{M}^{K}$, i.e., $\frac{1}{K}\sum_{i=1}^{K}\mathbb{E}[\mathcal{R}_{i}]$, where $\mathcal{R}_{i}$ denotes expected returns in MDP $M_{i}$. In order to train such policy, Wang et al., (2018) propose to use GNNs to model the relationships between different components of the state and action spaces across tasks, as GNNs are capable of capturing the structural and relational information inherent in various robot morphologies.
@@ 3 Entity-based Reinforcement Learning
To apply entity-based RL to Yawning Titan, we make use of the main constituents of the Entity Neural Network project of Winter et al., 2023a . That is, the Entity Gym interface (Winter et al., 2023b, ), the RogueNet policy architecture (Winter et al.,, 2021), and the Entity Neural Network Trainer framework for training agents in Entity Gym environments (Winter et al., 2023c, ). The following section is a summary of components of these works necessary for later discussion and doesn’t represent an original contribution of this paper. For more details on implementation and architecture, refer to the documentation of the repositories of Entity Gym (Winter et al., 2023b, ) and the Entity Neural Network Trainer (Winter et al., 2023c, ), as well as the blog post (Winter et al., 2023a, ) providing a comprehensive introduction to entity-based RL and the motivation behind the Entity Neural Network project.

### v3

@@ 1 Introduction
The development of autonomous agents for cyber defence is an area of research that aims to address the increasing complexity and volume of cyber threats. Solutions based on deep reinforcement learning (RL), inspired by high-profile successes in other complex domains (Vinyals et al.,, 2019; Roy et al.,, 2021; Bello et al.,, 2017), offer the potential to rapidly detect, analyse, and respond to cyber attacks at scale (Nguyen and Reddi,, 2021). A fundamental challenge involved in the deep RL approach is to develop agents that are generalisable to variable network topologies. This is difficult to achieve with fixed-input neural network architectures commonly used for function approximation and standard RL interfaces such as Gym or Gymnasium (Brockman et al.,, 2016; Towers et al.,, 2024), which expect fixed-size observation vectors and action spaces. The lack of this generalisability imposes serious limitations on the potential deployment of RL agents in real-world network scenarios, as enterprise networks can vary dramatically in size, structure, and complexity.
@@
This problem falls into the broad category of compositional generalisation, which is a long-studied problem in reinforcement learning (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Diuk et al.,, 2008). Recently, a number of works have explored introducing compositional generalisation to deep reinforcement learning  (Lin et al.,, 2023; Mambelli et al.,, 2022; Zhou et al.,, 2022; Haramati et al.,, 2024), through using permutation invariant policy architectures based on attention (Vaswani et al.,, 2017) or Deep Sets (Zaheer et al.,, 2017). Indeed, some existing work in autonomous cyber defence has suggested using Transformers to parameterise a value function (Mern et al.,, 2021), or using graph neural networks for policies (Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023) which are invariant to the number of nodes in the environment.
@@
This work aims to address this challenge by framing autonomous cyber defence as an entity-based reinforcement learning problem (Winter et al., 2023a, ). Entity-based RL is a paradigm where environments are decomposed into collections of discrete entities which an agent observes and acts upon directly. In the context of a network environment, each node of the network can be treated as an entity, with a defending agent’s observation space permitted to vary between environment instances according to how many nodes are visible. With a suitable training regime, it should be possible for an agent trained using entity-based RL to generalise to any network topology, provided the network in question contains entities of familiar types. The Entity Gym framework (Winter et al., 2023b, ), intended to occupy a similar role as OpenAI Gym or Gymnasium, provides a standard interface in which to cast entity-based environments, and allows any agent designed for entity-based reinforcement learning to be seamlessly deployed in such environments.
@@
In this paper, we introduce a wrapper for the Yawning Titan cyber-simulation environment (Andrew et al.,, 2022) that adheres to the Entity Gym interface. In this environment, we trained and compared two types of policies: a Transformer-based policy, RogueNet (Winter et al., 2023a, ), trained using the Entity Gym interface and a multilayer perceptron (MLP) policy trained on the standard Gym (Brockman et al.,, 2016) interface for Yawning Titan. We trained both types of policies using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017) with two different training modes, which we refer to as Random and Static, across network sizes of 10, 20, and 40 nodes. In the Random mode, a new random network topology is generated with each environment reset, whilst in the Static mode, a random network topology is initialised at the start of training and kept constant throughout. This approach has been designed to facilitate comparison with the MLP-based policy, which is able to train well in the Static mode, but struggles with Random training where it encounters a new network configuration with every episode. We also performed zero-shot evaluation of the Transformer-based policies on network sizes not seen during training, something not possible with standard MLP-based policies. From these experiments, we find that in this environment, entity-based policies outperform MLP-based policies during training, particularly in the Random mode where the MLP-based policy struggled to learn. We also find that entity-based policies generalise well to networks of unseen sizes, performing similarly to policies trained natively in those environments.
@@
The paper is organised as follows. In Section 2 we give an overview of motivating background and related literature. In Section 3 we describe the Entity Gym and the RogueNet architecture (Winter et al., 2023a, ; Winter et al., 2023b, ), which we have used to construct our environment and train entity-based policies. In Section 4 we provide a description of the Yawning Titan simulator and choices we have made for the environment setting for our experiments. In Section 5 we explain how we have adapted the Yawning Titan simulator to the entity-based framework. In Section 6 we outline the methodology and experiments we have chosen to run. In Section 7, we present the results of our experiments and conclude with a discussion of the results and future work in Section 8.
The code for our Entity Gym adaptation of Yawning Titan, as well as the scripts used to run the experiments in this paper, are available in the following repository: https://github.com/alan-turing-institute/Entity-Based-Yawning-Titan.
@@ 2.1 Autonomous Cyber Defence
The CybORG environment (Standen et al., 2022b, ) was initially conceived as part of a competition, Cage Challenge 1 (Standen et al., 2021a, ), intended to encourage research into RL-based autonomous cyber defence agentsngExamples of h solutions to the challenge include using agents trained with PPO (Schulman et al.,, 2017), combined with a bandit controller to select between policies specialised in specific attack patterns (Foley et al., 2022a, ; Hicks et al.,, 2023; Foley et al., 2022b, ; Bates et al.,, 2023), or with an ensemble approach (Wolk et al.,, 2022). Whilst not the focus of our paper, the CybORG environment features several types of nodes that could be used as entity types in the entity-based RL setting; User Host, Enterprise Server, Operational Host and the Operational Server.
Recognising that training and evaluating on a fixed network environment is not satisfactory for real-world applicability, a number of works have explored forms of generalisation. For example, to unseen red agent dynamics (Wolk et al.,, 2022), or to unseen network topologies, whether that be on fixed networks with variable connectivity (Collyer et al.,, 2022), or to networks containing differing numbers of nodes (Applebaum et al.,, 2022; Mern et al.,, 2021; Janisch et al., 2023a, ; Janisch et al., 2023b, ; Palmer et al.,, 2023). Indeed, the recent CAGE Challenge 4 (TTCP CAGE Working Group,, 2023) introduced a level of variability in the number of possible user nodes to encourage solutions addressing this problem.
@@
To the best of our knowledge, the most relevant prior work is Mern et al., (2021, 2022), where the authors implement a reinforcement learning environment for the autonomous defence of an industrial control network, and use an attention-based action-value function. Similarly to the entity-based paradigm, they use separate modules to encode different node-types, and share encoder parameters between all nodes of the same type. A global attention layer is then used to pass information between all node embeddings, before action-value estimates are decoded for all actions on all nodes. The parameters are learnt using a version of the Deep Q-Learning algorithm (Mnih et al.,, 2013; Hessel et al.,, 2017). This differs from the explicit policy parameterisation used in this work, trained using PPO. An advantage of a policy-based approach is that it allows for a more natural implementation of composite action spaces.
@@ 2.2 Entities and Objects in Reinforcement Learning
In this section we provide an overview of attempts to leverage the underlying structure of RL environments, as well as practical approaches for dealing with complex environments with potentially variable observation and action spaces, such as in the case of autonomous cyber defence. In the first instance, particularly in the case of factored MDPs (Boutilier et al.,, 2000), the motivation is to find more efficient solution methods by exploiting the ‘factorised’ structure of an environment. More broadly, all of these approaches contribute to solving the problem of compositional generalisation (Zhao et al.,, 2022; Lin et al.,, 2023). That is, training policies that are able to generalise across environments containing different compositions of familiar objects or entities and dynamics. For further reference, Mohan et al., (2024) provide a comprehensive survey of structure in reinforcement learning.
@@ 2.2.1 AlphaStar and entity-based Reinforcement learning
AlphaStar (Vinyals et al.,, 2019; Choi,, 2020) is an agent developed to play the game Starcraft II at human professional level. Starcraft II is a real-time strategy game in which two opposing players develop a base of operations and direct armies of units with the goal of defeating their opponent. Starcraft is a hard problem, both in terms of the practicalities of agent design and the complexity of strategies, which captures many of the challenges of real-world deployment of autonomous agents. Of particular relevance is that the number and nature of the entities an agent can observe and act upon varies throughout the game. This problem may be similar to that of nodes joining or leaving a network, or more generally the problem of transferring to an unfamiliar network consisting of familiar entities. Rather than using purely image-based inputs of the game screen as observations, AlphaStar maintains an entity list consisting of every visible object the agent can interact with. This could be army units, buildings, resources, or terrain. Crucially, this list will vary throughout a game, and between games.
Each entity is associated with a one-dimensional vector that summarises the state and properties of the entity. AlphaStar encodes these entity observations through a three-layer Transformer (Vaswani et al.,, 2017) to learn the relationships between entities. The entity embeddings generated by the Transformer are merged with other scalar and game-map information to form a joint representation for a single time step’s observation, which is used as input to an LSTM (Hochreiter and Schmidhuber,, 1997) that learns temporal dependencies.
@@ 2.2.2 Factored MDPs
Factored MDPs (Boutilier et al.,, 2000; Guestrin et al.,, 2003; Osband and Van Roy,, 2014) provide a framework that allows for more efficient learning on environments that may be naturally ‘factorised’ into distinct components. Here, the reward and transition functions can also be factorised according to these components, and display conditional independence properties. Guestrin et al., (2003) introduced efficient solution methods for reinforcement learning in factored MDPs as well as a toy environment, SysAdmin, on which to study them. In this environment, a system administrator must maintain a network of computers, each of which can be functional or failed. The administrator must decide whether to reboot one of the machines or do nothing at each time step, aiming to maximise rewards received for working machines, while considering that machine failures are influenced by the state of neighbouring machines. In this environment, the transition of each individual machine is conditionally independent of all other machines given its immediate neighbours, hence it can be cast as a factored MDP. The SysAdmin example bears resemblance to autonomous cyber defence environments, and the F-MDP framework could be a suitable candidate to express network security problems such as the ones considered in this work.
@@ 2.2.3 Relational Reinforcement Learning
Zambaldi et al., (2019) study ‘deep’ relational reinforcement learning, and advocate the use of GNN or attention-based function approximators in order to learn abstract representations of relations between entities in an environment. This is opposed to traditional relational RL, where relations are provided a-priori. They find the use of ‘relational’ attention modules allows an agent to generalise zero-shot to unseen environment configurations. This approach essentially discards the traditional formal solution methods based on logical descriptions of the environment. In other words, it is a purely entity-based approach, where the environment is split into distinct entities, and any relations between them are learnt during training.
@@ 2.2.5 Graph-based Reinforcement Learning for robotic control
A number of papers have studied the use of GNN or Transformer-based policies on robotics continuous control environments, which Hong et al., (2022) refers to as ‘Inhomogeneous Multitask Reinforcement Learning’. Here, the agent is tasked with learning policies across multiple tasks that may have different (inhomogeneous) state and action spaces or dynamics. The goal is to leverage shared knowledge and experience across these tasks to improve learning efficiency and performance. This has been successfully applied to robotics environments (Wang et al.,, 2018; Huang et al.,, 2020; Kurin et al.,, 2021; Hong et al.,, 2022), where the observation and action space are taken to be nodes in a graph of robot joints. An agent may be trained across different robot morphologies given by different graphs, with the goal that the policy is generalisable across all of these settings. Such tasks can be formalised by considering a set of MDPs $\mathcal{M}^{K}=\{M_{1},\dots,M_{K}\}$ (Kurin et al.,, 2021; Hong et al.,, 2022). The set $\mathcal{M}^{K}$ is inhomogeneous if there exists $i,j\in[K],i\neq j$ such that $dim(\mathcal{S}_{i})\neq dim(\mathcal{S}_{j})$ or $dim(\mathcal{A}_{i})\neq dim(\mathcal{A}_{j})$, where $\mathcal{S}_{i}$ and $\mathcal{A}_{i}$ refer to the state and action spaces of MDP $M_{i}$. The aim is then to construct a policy that maximises the average expected discounted return over all environments in $\mathcal{M}^{K}$, i.e., $\frac{1}{K}\sum_{i=1}^{K}\mathbb{E}[\mathcal{R}_{i}]$, where $\mathcal{R}_{i}$ denotes expected returns in MDP $M_{i}$. In order to train such a policy, Wang et al., (2018) propose using GNNs to model the relationships between different components of the state and action spaces across tasks.
@@ 3 Entity-based Reinforcement Learning
To apply entity-based RL to Yawning Titan, we make use of the main constituents of the Entity Neural Network project of Winter et al., 2023a . Namely, the Entity Gym interface (Winter et al., 2023b, ), the RogueNet policy architecture (Winter et al.,, 2021), and the Entity Neural Network Trainer framework for training agents in Entity Gym environments (Winter et al., 2023c, ). The following section is a summary of components of these works necessary for later discussion and does not represent an original contribution of this paper. For more details on implementation and architecture, refer to the documentation of the repositories of Entity Gym (Winter et al., 2023b, ) and the Entity Neural Network Trainer (Winter et al., 2023c, ), as well as the blog post (Winter et al., 2023a, ) providing a comprehensive introduction to entity-based RL and the motivation behind the Entity Neural Network project.